# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v2

@@ 3.1 Spatial Discretization Scheme
Figure 2: Iterations for finding the supporting baseline without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum. The solid line is the original spectrum and the dotted line is the evaluated supporting baseline. With smooth extrapolation to the original spectrum, the speed of convergence has significant improvement.
@@ 3.2 Treatment of Boundary Conditions
The primal problem requires the supporting baseline $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$. For low and high pass signals of which the spectrum value are high and drop abruptly at either of the ends, this will make $g(x)$ inherits the trend of $f(x)$ at both ends at some interval so that convergence in this case will be very slow with reasonable grid space. To alleviate this effect, we adopt a smooth extrapolation strategy by symmetrically appending gently decaying segments at both ends to make the extended value align with the minimum of $f(x)$, and the extended derivative smoothly approach to zero within the extended region. This strategy significantly improves the condition number of the coefficient matrix of the discrete system, ensuring the numerical stability of the iterative solution, without any altering the primal spectrum within the original interval. Fig. 2 shows the illustrations.
3.3 Postprocessing: From Baseline to Mode Identification
Upon the above course when the algorithm converges, we obtain the optimal discrete baseline vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly. However, there are some residual background fluctuations at the bottom that hinders one to objectively distinguish those modes. To solve this problem, we assume that the residual consists of two components: first, a background fluttering component originate from noise or minor undulations, of which the amplitude concentrates within a very small range; second, significant protrusions from original modes, which are larger compared with the former and more scattered. In this view, when investigating the value of the residual spectrum value distribution, we can intuitively find that most values concentrate near the level where fluttering component resides. This inspires us to employ Kernel Density Estimation (KDE) to non-parametrically model the spectrum value distribution as a probability density function:
@@ 3.3 Postprocessing: From Baseline to Mode Identification
Figure 3: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function. Fig. 3a shows the converged $g(x)$ and Fig. 3b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 3c shows the estimate kernel density pertaining the gap at the bottom. 3d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
@@ 3.4 The Whole Workflow
Based on the above discrete framework, each step of Algorithm 1 can be implemented by specific linear algebraic operations. The whole workflow can be summarized as Algorithm 2. We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the supporting baseline function $g(x)$. After that we smoothly extend $f(x)$ at both ends and begin Algorithm 1 until it converges. Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, so as to isolate band components.
@@
10: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
11: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
12: Output $\{u_{k}\}$
4 Numerical Results
@@
All experiments are conducted under the same parameter settings unless otherwise stated. The stopping criterion is set as $\dfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$. The number of discrete grid points is fixed at $N=200$ for the normalized frequency domain $[0,1]$. Detailed descriptions of the test signals and comparative protocols are provided in the corresponding subsections.
Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
4.1 Experiments on Finding the Supporting Baseline
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. We only illustrate the initial and final state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in C.
@@ 4.1 Experiments on Finding the Supporting Baseline
Experiment 2: The second example is a multi-mode signal, $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$, which has two modes, exactly. The normalized center frequency is 0.05 and 0.1, which is relatively very close to each other. However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the support baseline forms a small bulge at the close frequency points, which just separates the two frequency points.
Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to time series (62). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
@@
Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
@@
Figure 8: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
@@ 4.2 Experiments on Central Frequencies
Table 1: The evaluated center frequencies
Experiment Ours Original VMD Max Difference
1 0.0550 0.0496 0.0054
2 0.0300, 0.0575 0.0251, 0.0507 0.0068
3 0.0100, 0.0775 0.1550, 0.2075 0.0008, 0.0530 0.1510, 0.2 0.0245
4 0.0125, 0.0875 0.2475 6.46e-5, 0.0067 0.0803 0.1682
5 0.005,0.0775 0.1, 0.125 0.1550, 0.1800 0.2075, 0.2325 0.2575, 0.2825 0.3075 0.000845, 0.0747 0.0999, 0.1250 0.1510, 0.182 0.223, 0.25 0.2760, 0.303 0.339 0.0315
Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s. As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD. It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.7c even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
@@ 4.3 Experiments Compared with Variational Mode Decomposition
In this section, we pick 2 typical signals for illustration. The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the supporting baseline. For the completeness of our comparison, we investigated all stopping criterion implemented in the source code of SVMD. The method considering about eliminating noise(northwest) drops the base-frequency and a batch of high frequency, while method for extract reconstruction, method based on Beyesian Estimation, or method considering the power of last mode all result in some redundant modes with low energy on the decomposed spectrum (see colored spectrum at the bottom in each subplot), which are not the essential modes that should included in the decomposition result (Fig. 11). In contrast, our method(Fig. 10) accurately extracts all necessary modalities without any redundant ones.
@@
Figure 14: The decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering noise(northwest), Bayesian Estimation(southwest), power of the last mode(southeast). Note that the figure for the exact decomposition is missed since the SVMD failed to converge in 10 seconds.
@@ 4.4 Experiment on Real Signals
Figure 15: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to MLII lead from the 100th record.
Figure 16: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to V4 lead from the 124th record.
In this section we investigate the performance of our method on real signals. As ECG signals are typical ones that combined with several narrow-banded components, we choose ECG signals as real signals for verification. Our ECG signals derives from MIT-BIH Arrhythmia Database [23, 24], and we choose randomly two signals for analysis, wherein the first is the MLII lead from the 100th record and the second is the V4 lead from the 124th record for our experiments respectively. Since each original signal has about 65000 samples which makes redundant for our experiment, we choose the first 2000 samples and do down-sampling at the ratio 1:10 for computation efficiency. For the limitation of the page we put the course of iteration and the decomposition result in Supplementary Materials(Fig. S3.16, Fig. S3.17 and Fig. S3.18, Fig. S3.19 for iteration and decomposition result with respect to the first and second chosen signals respectively), and instead we analyse the correlation coefficient(CC, the higher the better) and the ratio between the power ratio(PR) with respect to the residual and the original signal(the lower the better), and illustrate the gap between the reconstructed signal from the sum of the IMFs that acquired by our algorithm and the source signal.
For real signals composed of narrow-banded components, our algorithm can obtain effective components, since on one side the gap between the source signal(solid line in Fig. 15 and Fig. 16) and reconstructed signal(dashed line in Fig. 15 and Fig. 16) is typically small, meaning that one can always revive the source signal with small error by adding all the IMFs obtained from algorithm. On the other side, as shown in Table 2, the sum of IMF components obtained by our algorithm can achieve high correlation coefficient(both over 0.99) and low power ratio with respect to residual and the source signal(at the order of $10^{-4}$), which indicates our algorithm can automatically retrieve almost all the meaningful components to the source signal and seldom ignore the significance of components with small energy.
Table 2: The number of IMFs acquired from our algorithm, the CC(Correlation Coefficients between the reconstructed signal and source signal) and the PR(Power Ratio with respect to the residual and the source signal) result.
Experiment Number of IMFs $C$ PR
MLII lead from the 100th record 15 0.9989 4.2635e-04
V4 lead from the 124th record 14 0.9959 2.0626e-04

### v3

@@ 3.1 Spatial Discretization Scheme
Figure 2: Iterations for finding the Cutting Curve without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum. The solid line is the original spectrum and the dotted line is the evaluated Cutting Curve. With smooth extrapolation to the original spectrum, the speed of convergence has significant improvement.
@@ 3.2 Treatment of Boundary Conditions
The primal problem requires the Cutting Curve $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$. For low and high pass signals of which the spectrum value are high and drop abruptly at either of the ends, this will make $g(x)$ inherits the trend of $f(x)$ at both ends at some interval so that convergence in this case will be very slow with reasonable grid space. To alleviate this effect, we adopt a smooth extrapolation strategy by symmetrically appending gently decaying segments at both ends to make the extended value align with the minimum of $f(x)$, and the extended derivative smoothly approach to zero within the extended region. This strategy significantly improves the condition number of the coefficient matrix of the discrete system, ensuring the numerical stability of the iterative solution, without any altering the primal spectrum within the original interval. Fig. 2 shows the illustrations.
3.3 Postprocessing: From Cutting Curve to Mode Identification
Upon the above course when the algorithm converges, we obtain the optimal discrete vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly. However, there are some residual background fluctuations at the bottom that hinders one to objectively distinguish those modes. To solve this problem, we assume that the residual consists of two components: first, a background fluttering component originate from noise or minor undulations, of which the amplitude concentrates within a very small range; second, significant protrusions from original modes, which are larger compared with the former and more scattered. In this view, when investigating the value of the residual spectrum value distribution, we can intuitively find that most values concentrate near the level where fluttering component resides. This inspires us to employ Kernel Density Estimation (KDE) to non-parametrically model the spectrum value distribution as a probability density function:
@@ 3.3 Postprocessing: From Cutting Curve to Mode Identification
After that, we use DBSCAN to merge adjacent peaks with low energy at the bottom for $f(x)-g^{*}(x)-t$.
Figure 3: Illustration of the non-consistence of the error between the converged optimal Cutting Curve and the original function. Fig. 3a shows the converged $g(x)$ and Fig. 3b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 3c shows the estimate kernel density pertaining the gap at the bottom. 3d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
@@ 3.4 The Whole Workflow
Based on the above discrete framework, each step of Algorithm 1 can be implemented by specific linear algebraic operations. The whole workflow can be summarized as Algorithm 2. We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the Cutting Curve function $g(x)$. After that we smoothly extend $f(x)$ at both ends and begin Algorithm 1 until it converges. Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, and aggregate adjacent minor peaks into single coherent intrinsic mode by DBSCAN.
@@
10: Use DBSCAN to merge spatially adjacent minor peaks into a single coherent intrinsic mode.
11: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
12: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
13: Output $\{u_{k}\}$
4 Experiments
@@
All experiments are conducted under the same parameter settings unless otherwise stated. The stopping criterion is set as $\dfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$. To adaptively determine the effective frequency band, we compute the cumulative energy distribution of the single-sided power spectrum. The cut-off frequency $f_{c}$ is defined such that
(64) $\cfrac{\int_{0}^{f_{c}}S(f)\text{d}f}{\int_{0}^{+\infty}S(f)\text{d}f}=0.95,$
which is found by scanning from the Nyquist frequency downward until the energy condition is met. The resulting $f_{c}$ effectively isolates the dominant spectral components while discarding high-frequency noise tails. The truncated spectrum is then down‑sampled to 100 points, and the optimal cutting curve is determined on this down‑sampled spectrum.
Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum Cutting Curve of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
4.1 Experiments on Finding the Cutting Curve
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. We only illustrate the initial and final state of the iteration for the evaluation of the cutting curve, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in Appendix C.
@@ 4.1 Experiments on Finding the Cutting Curve
Experiment 2: The second example is a multi-mode signal, $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$, which has two modes, exactly. The normalized center frequency is 0.05 and 0.1, which is relatively very close to each other. However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the cutting curve forms a small bulge at the close frequency points, which just separates the two frequency points.
Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the Cutting Curve of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the Cutting Curve of FFT with respect to time series (65). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed Cutting Curve precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
@@
Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the cutting curve of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
Figure 8: to be continued
@@
Figure 9: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the cutting curve of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
@@ 4.2 Experiments on Central Frequencies
Figure 10: The theoretical frequencies (solid line), our estimated frequencies (dashed line) and the VMD-evaluated frequencies (dash-dotted line) after iteration with respect to Experiment 1-5 (a)-(e).
Fig. 10 shows the theoretical center frequencies, our estimated frequencies and VMD-evaluated frequencies with respect to the experimental synthetic signals in order, where the solid line stands for the theoretical frequency of each mode, the dashed line stands for our estimation and the dotted-dash line stands for the ones after VMD’s evaluation, and different colors stands for different decomposed modalities. It can be seen that our estimated values are very close to the actual modal frequencies, and the accuracy of the estimation is independent of the number of modes to be decomposed, although for VMD procedure there are cases in which the central frequencies of modes with small energy will deviate from the theoretical.
@@ 4.3 Experiments Compared with Variational Mode Decomposition
In this section, we pick 2 typical signals for illustration. The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the Cutting Curve. For the completeness of our comparison, we investigated all stopping criterion implemented in the source code of SVMD. The method considering about eliminating noise(northwest) drops the base-frequency and a batch of high frequency, while method for extract reconstruction, method based on Beyesian Estimation, or method considering the power of last mode all result in some redundant modes with low energy on the decomposed spectrum (see colored spectrum at the bottom in each subplot), which are not the essential modes that should included in the decomposition result (Fig. 13). In contrast, our method(Fig. 12) accurately extracts all necessary modalities without any redundant ones.
@@
Figure 16: Local zoom of the decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering Bayesian Estimation(left), power of the last mode(right). Note that the figure for the one considering noise and the one with exact decomposition is missed since the SVMD failed to converge in 10 seconds.
@@ 4.4 Experiment on Real Signals
To demonstrate the generality for different application domains, we also evaluate its performance on electrocardiogram (ECG) signals. ECG signals are representative examples of real-world non-stationary physiological signals composed of multiple characteristic waveforms (such as the P wave, QRS complex, and T wave), each occupying distinct frequency bands. Accurately decomposing an ECG signal into its constituent modes is therefore a challenging yet practically meaningful test for any adaptive mode decomposition framework.
We apply our method to selected records from the MIT-BIH Arrhythmia Database. Specifically, we extract 2000-sample segments from the MLII lead of Record 100, the V4 lead of Record 124, and the V2 lead of Record 102. For each signal, our method automatically determines the number of intrinsic mode functions and their corresponding initial center frequencies, which are then applied to VMD for the final decomposition.
Fig.17-Fig.19 shows the decomposition result of MLII lead of Record 100, the V4 lead of Record 124, and the V2 lead of Record 102. In each image, the top subfigure represents the original signal, the bottom subfigure represents the residual signal, and the middle graphs represent the decomposed IMFs. (b) stands for the spectrums of each decomposed IMF, (c) shows the final converged cutting curve for the spectral amplitude of each signal, and (d) shows the reconstructed signal (orange) with respect to the source signal (blue). It is shown that our algorithm determines 4, 4, 2 modes for these three signals in order. At the time domain we can see that in each decomposed IMFs in our chosen number of mode can reflect the basic characteristic that is helpful for the disease diagnosis. For example, in Fig.17(a) we see the heartbeat rhythm from IMF 1-3 and IMF 4 amplifies the characteristic before and after the peaks. The similar phenomenon also happens in Fig.18(a). In Fig.19We see that each modality maintains the significant shape features of the original signal. The Correlation Coefficients between the reconstructed and source signal is 0.9993, 0.9986, 0.9998 respectively.
Fig.20 shows the performance comparison between our preferred mode number and center frequencies with respect to randomly chosen mode number from 2 to 10 for each investigated signal. The first line is the result for MLII Lead of Record 100, the second line is the result for V4 Lead of Record 124 and the third line is the result for V2 Lead of Record 102. (a),(c),(g) stands for the performance of orthogonality (average value of off-diagonal elements), (b),(e),(h) stands for the Correlation Coefficients performance and (c),(f),(i) stands for Power Ratio of Residual performance. The red points stands for the performance under our chosen mode number and center frequencies, while the blue lines stands for the performance from randomly choosing mode number from 2 to 10. The first 2 lines shows that our algorithm can choose a reasonable mode number protecting over-decomposition, while the third line shows that the V2 Lead of Record 102 seems under-decomposition.
However, Fig.21-Fig.23 shows that by artificially increasing the mode number it seems increasing some redundant mode or similar mode. In Fig.21 and Fig.22 the first 5 modes are similar in that they all reflect the heart beat rhythm, although they are almost orthogonal due to their frequency increasing relationship with each other. However, this does not bring any improvement to the interpretability for the help of disease diagnosis. Especially, in Fig.23 we see when setting mode number into 3, the decomposed modes can hardly reflect the basic characteristic of the source signal. In contrast, the performance of orthogonality in turn improves with the increasing of the mode number as in Fig.21 and Fig.22 the former 5 modes are almost orthogonal due to their frequency increasing relationship with each other. This indicates that evaluating the quality of signal decomposition cannot be solely based on performance indicators, but also on the physical meaning of the actual decomposed signal. To some extent, our modal determination method is naturally related to the geometric shape of the signal spectrum, and in most cases, the number of modes is determined based on the identifiability of the spectral peaks. The resulting modes can be proven valuable in engineering scenarios.
Figure 17: The Decomposition result for MLII Lead of Record 100. (a) The original signal(top), the decomposed IMFs(middle 4), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cutting line, (d) The source signal and the reconstruction signal.
Figure 18: The Decomposition result for V4 Lead of Record 124. (a) The original signal(top), the decomposed IMFs(middle 4), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cuttnig curve, (d) The source signal and the reconstruction signal.
Figure 19: The Decomposition result for V2 Lead of Record 102. (a) The original signal(top), the decomposed IMFs(middle 2), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cutting curve, (d) The source signal and the reconstruction signal.
Figure 20: The performance of decomposition of our chosen K and center frequencies for each ECG signal compared with random select K from 2 to 10 to proceed VMD.
Figure 21: The decomposition result on MLII Lead for Record 100 when K is set over our estimation (for example 7).
Figure 22: The decomposition result on V4 Lead for Record 124 when K is set over our estimation (for example 7).
Figure 23: The decomposition result on V2 Lead for Record 102 when K is set over our estimation (for example 3).